OpenStudy (anonymous):

Find a number t such that the distance between (-2,1) and (3t, 2t) is as small as possible.

OpenStudy (bahrom7893):

What's the distance formula?

OpenStudy (bahrom7893):

d = sqrt( (x2-x1)^2 + (y2-y1)^2 )

OpenStudy (bahrom7893):

d = sqrt( (3t+2)^2 + (2t-1)^2 )

OpenStudy (bahrom7893):

d = sqrt( (9t^2+12t+4) + (4t^2-4t+1) ) d = sqrt( 9t^2 + 4t^2 + 12t - 4t + 4 + 1 )

OpenStudy (bahrom7893):

d = sqrt (13t^2 + 8t + 5) d' = (1/2) (13t^2 + 8t + 5)^(-1/2)*(26t+8)

OpenStudy (bahrom7893):

So now solve: 26t+8 = 0 and 13t^2 + 8t + 5 = 0 Get your ts and plug them all into d = sqrt( (3t+2)^2 + (2t-1)^2 ). See at which t d is smallest, and that will be your t.